OpenStudy (anonymous):

y=mx+b so how does that work with x+ 2y=7?

OpenStudy (anonymous):

solve for y 2y = 7 - x y = (7 - x)/2

OpenStudy (anonymous):

^ Correct: y = -1/2(x) +7/2 Therefore, m = -1/2 b = 7/2

OpenStudy (anonymous):

... y = 7/2 - x/2 y = -x/2 - 7/2

OpenStudy (anonymous):

did that help, help needed?

OpenStudy (anonymous):

Oiler, it's 7/2, not -7/2

OpenStudy (anonymous):

Yes, it's what we put for the equation, but we don't understand how to graph that and don't know how to use the online grapher we looked at

OpenStudy (anonymous):

of course, y = 7/2 - x/2 y = -x/2 + 7/2

OpenStudy (anonymous):

Here's a way to graph the original form: ...

OpenStudy (anonymous):

Ok, we graphed that line, but it's not intersecting with the other equation 3x+y =6

OpenStudy (anonymous):

put in zero for x, 0 + 2y = 7, y = 7/2, gives the point (0,3.5) put in zero for y, x + 0 = 7, x = 7, gives the point (7,0) Real easy, without solving for y

OpenStudy (anonymous):

For 3x+y =6 put in zero for x, 0 + y = 6, y = 6, gives the point (0,6) put in zero for y, 3x + 0 = 6, x = 2, gives the point (2,0) Real easy, without solving for y

OpenStudy (anonymous):

I'm trying it out

OpenStudy (anonymous):

You're getting a medal, it worked!

OpenStudy (anonymous):

Nice job oiler :D

OpenStudy (anonymous):

You too! Thanks so much.

OpenStudy (anonymous):

Oiler are you still there?

OpenStudy (anonymous):

2x+3y=11 (0, 11/3) 3x-2y=_16 (0,8) how do I graph 11/3?